# Comparison of bar formation mechanisms II: does a tidally induced bar grow faster than an internally developed bar?

### v1

@@ 4.1 Do tidal bars grow faster?
Our results suggest the fundamental mechanism driving bar growth in cold and warm disks may still be their internal bar instability even under tidal perturbations. The external perturbation mainly affects the onset of the bar formation. Bars formed under tidal interactions in these disks are not purely tidally induced but rather a result of both internal instability and external perturbation. For convenience, we continue to use the term “tidally induced bars” loosely throughout this paper.
@@
Note that $\tau_{\mathrm{bar}}$ is not precisely identical between tidally induced bars and internally developed ones. We propose two potential factors for this minor discrepancy. One is the transient spiral arms that contribute to the $A_{2}$ signal with the extreme cases being those of the prograde interactions with a mass ratio of 1/1. The other factor is the effect of the DM halo spin. Although the primary galaxy’s halo is initially set up without net angular momentum, the tidal force from the perturber spins it up. Isolated simulations have demonstrated the significance of inner halo angular momentum in bar formation: bars grow more rapidly in prograde halos while more slowly in counter-rotating halos (see Figure 3. in Kataria & Shen, 2022). Observations also suggest a relationship between bar structures, dark matter halo mass/spin, and galaxy mass/spin (Romeo et al., 2023). The counter-rotating halo spin may explain why the retrograde 1/1 interaction of the cold disk has a larger $\tau_{\mathrm{bar}}$.
Although likely a secondary factor, the spinned-up halo does contribute to bar formation. A detailed examination of the halo spin effect in interactions will be addressed in subsequent papers of this series.
@@ 4.2 How perturbation affects the onset of the bar formation
We analyze the flyby interaction of the warm disk model under an intermediate perturbation ($M_{\mathrm{per}}/M_{\mathrm{gal}}=1/3$, perpendicular orbit), in which the bar formation is advanced by 0.7 Gyr. The onset of the large-scale bar-like perturbation in this case is presented in Figure 4. The left column shows the stellar surface density maps. The right three columns illustrate the deviations between the current snapshot and the initial condition, specifically in terms of stellar surface density ($\Delta\Sigma$), averaged radial velocity ($\Delta\langle V_{R}\rangle$), and averaged azimuthal velocity ($\Delta\langle V_{\phi}\rangle$), respectively. We note that the average radial velocity of the axisymmetric disk at the beginning of the simulation is zero Thus, $\Delta\langle V_{R}\rangle=\langle V_{R}\rangle$.
The first row depicts a time shortly after the perturber’s closest approach. The interaction triggers a perturbation wave that propagates to the galaxy’s center. This wave generates a “seed bar”, as indicated by the quadrupole signal in the velocity fields shown in the second row. Subsequently, the swing amplification feedback loop amplifies this seed bar into a robust bar structure. We illustrate the bar’s evolution at $t_{0}$, i.e. the time when $A_{2}$ reaches 0.1, in the third row. The final row displays the bar at a later stage when $A_{2}$ reaches 0.2, a common criterion for identifying strong bars.
To further confirm the link between the perturbation and the onset of the bar formation, we rerun several flyby simulations of the warm disk with pericenter time being delayed by 1 Gyr. As expected, the bar formation is delayed by a similar amount compared with the original runs as revealed by the $t_{0}$ parameter.
@@
For comparison, we present the onset of bar formation in the same disk evolved in isolation in Figure 5. A similar quadrupole signal in the velocity fields spontaneously emerges at the galaxy center as shown in the first row. However, the “seed bar” appears 0.7 Gyr later compared to the flyby interaction scenario shown in Figure 4. The bottom three rows of Figure 5 illustrate the development of the “seed bar” into a strong bar structure. The evolution of the bar in the isolated case is similar to that under tidal perturbation (Figure 4) except the stronger spiral arms in the latter case. This consistent evolution further supports the notion that the tidal bar grows at a rate comparable to its spontaneous counterpart in the warm disk.
@@
Figure 6: Same as Figure 4 but for a delayed case (cold disk, 1/3, perpendicular). Although weak, the cold disk spontaneously generates a “seed bar” prior to the interaction (top row). When the perturbation wave reaches the galaxy’s center, it is out of phase with the pre-existing seed bar, causing destructive interference and limiting the bar’s growth (middle two rows). Despite this initial delay, the bar eventually resumes steady growth at a later stage (bottom row).
@@
Tidal perturbations can also postpone the onset of bar formation. Figure 6 illustrates a delayed scenario in the cold disk model ($M_{\mathrm{per}}/M_{\mathrm{gal}}=1/3$, perpendicular orbit). The first row shows a moment shortly before the perturber’s closest approach. Although weak, the cold disk has already spontaneously generated a “seed bar” prior to the interaction. When the perturbation wave reaches the galaxy center, it interferes destructively with the original seed bar and limits the bar’s growth. Ultimately, one mode prevails over the other at a later time and starts the steady growth of the bar as seen in the bottom row ($A_{2}=0.1$). We propose that the internal mode dominates over the external mode as the perturber is already more than $300\;R_{d}$ away when $A_{2}$ reaches 0.1.
@@
Our results align with the interpretation of Moetazedian et al. (2017) regarding the delayed case. Destructive interference is due to the external perturbation wave being accidentally out of phase with the internally developed seed bar when it reaches the galaxy center. We test this hypothesis by re-simulating the delayed case with slightly altered pericenter times, shifted by ±$\pm 0.1$ Gyr. In both the earlier and later interactions, the onset of bar formation is advanced by approximately 0.2 Gyr compared to the isolated cases, rather than being delayed as in the original interaction. This result underscores the difficulty for tidal perturbations to delay bar formation within galaxies inherently susceptible to internal instability, thereby explaining why the majority of our interaction simulations have earlier emergences of bars. Tidal perturbations provide stronger initial noises for the swing amplification feedback loop, while the probability of the external perturbation wave being out of phase with the internally developed seed bar is indeed small. Our results also highlight the complexity of predicting the galaxy morphology which is influenced by both their environment and internal baryonic physics (Zhou et al., 2020).
4.3 bar growth in hot disk model
@@
In a more specific context, the term “tidally induced bars” refers exclusively to bars in galaxies that are stable against bar formation when isolated. Consequently, we examined bar growth in the hot disk model, which avoids bar instability throughout a 6 Gyr isolated evolution.
@@
We utilize the same exponential fitting (Equation 3) for the $A_{2}$ evolution in the hot disk model. Since $A_{2}$ does not exhibit a distinct peak in the hot disk, we extend the upper limit of the fitting time range to the end of the simulation, i.e., 6 Gyr. For the lower limit, we set $A_{2}>0.025$ and exclude the initial 0.75 Gyr of the simulation. Visual examination of the stellar surface density indicates that the significant increase in the $A_{2}$ signal around the interaction time ($t_{\rm peri}=0.5$ Gyr) is primarily due to overall distortion of the disk and transient spirals, rather than bar formation. The fitting outcomes are presented in Figure 7. Unlike the cold and warm disks, the hot disk does not display a consistent $\tau_{\mathrm{bar}}$, ranging from a minimum of 4.20 Gyr to a maximum of 12.81 Gyr. Notably, the $M_{\mathrm{per}}/M_{\mathrm{gal}}=1/1$ interactions yield larger $\tau_{\mathrm{bar}}$values compared to the weaker interactions even though the bars in the former cases are stronger.
@@
where $k$ represents the linear growth rate of the bar strength. The linear fitting is also presented in Figure 7 with the $k$ values noted in the text box. We find that stronger interactions correspond to larger $k$ values, suggesting the correlation between interaction strength and bar growth in the hot disk.
Visually, the linear function seems to offer a slightly better fit to the bar growth in the hot disk for most scenarios. From a statistical perspective, the linear fit yields a smaller reduced chi-squared statistic ($\chi_{\nu}^{2}$) compared to the exponential fit. Conversely, the cold and warm disks exhibit a more favorable fit with the exponential function during the bar formation phase. The linear fitting approach may be more appropriate for describing the secular evolution of the bar in the cold and warm disks, occurring after the buckling stage and prior to the saturation stage (refer to the inset axes in Figure 1, and also Figure 4 in Paper I).
The tidally induced bars in the hot disk deviate from the exponential growth profile, with their linear growth rate being linked to the interaction strength. This contrasts with the cold and warm disks, where bar growth is primarily driven by internal disk instability. This difference makes it less meaningful to compare the growth timescale $\tau_{\mathrm{bar}}$ between the hot disk and the cold/warm disks. It also suggests that bars in the hot disk may not operate under the same growth mechanism as those in the cold and warm disks. Possibly, the hot dynamical state of the disk resists the disturbance caused by the perturber, preventing the disruption of the galaxy center and rendering it less susceptible to the swing amplification feedback loop (Dubinski et al., 2008).
@@
If the term “tidally induced bars” is strictly applied to those in galaxies that inherently avoid bar instability, their formation and evolution differ from those in galaxies that are prone to bar instability. Quantitative comparison of bar shapes and dynamical/kinematical analysis of the bar regions have the potential to distinguish these two categories of bars. This will constitute the long-term objective of this series of papers.
@@ 5 Summary
We apply an exponential fitting to the $A_{2}$ evolution in cold and warm disks to quantify the growth timescale $\tau_{\mathrm{bar}}$ and the onset time $t_{0}$ of the bar formation (Figure 1). The tidally induced bars show similar $\tau_{\mathrm{bar}}$ to that of the internally developed bars with differences less than 20%, indicating that tidal bars have similar growth rates to their spontaneous counterpart in the same disk (Figure 2). The tidal perturbation promotes/delays the bar formation by advancing/postponing the onset time $t_{0}$. This conclusion is further corroborated by comparing the $A_{2}$ evolution against $(t-t_{0,\ fit})$ (Figure 3). If the disk inherently possesses instability, the primary driver of bar formation is its internal nature that remains largely unaffected by tidal perturbations.
We further investigate the onset of the bar formation by examining the perturbation wave. In instances where the perturbation advances bar formation, the perturbation wave propagates the galaxy center shortly after the perturber’s pericenter and evokes a “seed bar” (Figure 4). The isolated case shows a similar evolution of the “seed bar” but its spontaneous emergence occurs approximately 0.7 Gyr later than in the interaction case (Figure 5). We also show a delayed interaction case where the perturbation wave is out of phase with the internally developed seed bar, leading to destructive interference and limiting the growth of the bar (Figure 6).
In the hot disk model, the tidally induced bars do not show consistent $\tau_{\mathrm{bar}}$ values (Figure 7) probably due to the deviation from the exponential growth profile. The linear fitting provides a slightly better description of the bar growth in the hot disk than the exponential fitting. This fact makes it less meaningful to compare the growth timescale $\tau_{\mathrm{bar}}$ between the hot disk and the cold/warm disks and hints that bars in the hot disk may not follow the same growth mechanism as those in the cold and warm disks.
However, the difference in the bar growth mechanisms holds the potential to separate the strictly-speaking “tidally induced bars” from those developed in galaxies that are inherently susceptible to bar instability. Further investigation of the bar’s shape and dynamical/kinematical analysis of the bar regions could be important in distinguishing between these two types of bars.
@@ Acknowledgements
We thank the anonymous referee for suggestions that helped to improve the presentation of the paper. We thank Zhi Li and Sandeep Kumar Kataria for their valuable insights on simulations and analysis. We also thank Rui Guo for their helpful discussions. The research presented here is partially supported by the National Key R&D Program of China under grant No. 2018YFA0404501; by the National Natural Science Foundation of China under grant Nos. 12025302, 11773052, 11761131016; by the “111” Project of the Ministry of Education of China under grant No. B20019; and by the Chinese Space Station Telescope project. J.S. acknowledges support from the Newton Advanced Fellowship awarded by the Royal Society and the Newton Fund. X.W. wishes to thank the financial support from the Natural Science Foundation of China (Numbers NSFC-12073026, NSFC-12433002). This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.

### v3

@@ 4.1 Do tidal bars grow faster?
Our results suggest the fundamental mechanism driving bar growth in cold and warm disks may still be their internal bar instability even under tidal perturbations. External perturbation mainly affect the onset of the bar formation. Bars formed under tidal interactions in these disks are not purely tidally induced but rather a result of both internal instability and external perturbation. For convenience, we continue to use the term “tidally induced bars” loosely throughout this paper.
@@
Note that $\tau_{\mathrm{bar}}$ is not precisely identical between tidally induced bars and internally developed ones. We propose two potential factors for this minor discrepancy. One is the transient spiral arms that contribute to the $A_{2}$ signal with the extreme cases being those of the prograde interactions with a mass ratio of 1/1. The other factor is the effect of the DM halo spin. Although the primary galaxy’s halo is initially set up without net angular momentum, the tidal force from the perturber spins it up. Isolated simulations have demonstrated the significance of inner halo angular momentum in bar formation: bars grow more rapidly in prograde halos while more slowly in counterrotating halos (see e.g. Collier & Madigan, 2021; Kataria & Shen, 2022; Joshi & Widrow, 2024). Observations also suggest a relationship between bar structures, DM halo mass and spin, and galaxy mass and spin (Romeo et al., 2023). The counterrotating halo spin may explain why the retrograde 1/1 interaction of the cold disk has a larger $\tau_{\mathrm{bar}}$.
Although likely a secondary factor, the spun-up halo does contribute to bar formation. A detailed examination of the halo spin effect in interactions will be addressed in subsequent papers of this series.
@@ 4.2 How perturbation affects the onset of the bar formation
We analyze the flyby interaction of the warm disk model under an intermediate perturbation ($M_{\mathrm{per}}/M_{\mathrm{gal}}=1/3$, perpendicular orbit), in which the bar formation is advanced by 0.7 Gyr. The onset of the large-scale bar-like perturbation in this case is presented in Figure 4. The left column shows stellar surface density maps. The right three columns illustrate the deviations between the current snapshot and the initial condition, specifically in terms of stellar surface density ($\Delta\Sigma$), averaged radial velocity ($\Delta\langle V_{R}\rangle$), and averaged azimuthal velocity ($\Delta\langle V_{\phi}\rangle$), respectively. We note that the average radial velocity of the axisymmetric disk at the beginning of the simulation is zero. Thus, $\Delta\langle V_{R}\rangle=\langle V_{R}\rangle$.
The first row depicts a time shortly after the perturber’s closest approach. The interaction triggers a perturbation wave that propagates to the galaxy’s center. This wave generates a “seed bar”, as indicated by the quadrupole signal in the velocity fields shown in the second row. Subsequently, the swing amplification feedback loop amplifies this seed bar into a robust bar structure. We illustrate the bar’s evolution at $t_{0}$, i.e., the time when $A_{2}$ reaches 0.1, in the third row. The final row displays the bar at a later stage when $A_{2}$ reaches 0.2, a common criterion for identifying strong bars.
To further confirm the link between the perturbation and the onset of the bar formation, we rerun several flyby simulations of the warm disk with the pericenter time being delayed by 1 Gyr. As expected, the bar formation is delayed by a similar amount compared with the original runs, as revealed by the $t_{0}$ parameter.
@@
For comparison, we present the onset of bar formation in the same disk evolved in isolation in Figure 5. A similar quadrupole signal in the velocity fields spontaneously emerges at the galaxy center as shown in the first row. However, the “seed bar” appears 0.7 Gyr later compared to the flyby interaction scenario shown in Figure 4. The bottom three rows of Figure 5 illustrate the development of the “seed bar” into a strong bar structure. The evolution of the bar in the isolated case is similar to that under tidal perturbation (Figure 4) except for the stronger spiral arms in the latter case. This consistent evolution further supports the notion that the tidal bar grows at a rate comparable to its spontaneous counterpart in the warm disk.
@@
Figure 6: Same as Figure 4 but for a delayed case (cold disk, 1/3, perpendicular). Although weak, the cold disk spontaneously generates a “seed bar” prior to the interaction (top row). When the perturbation wave reaches the galaxy’s center, it is out of phase with the preexisting seed bar, causing destructive interference and limiting the bar’s growth (middle two rows). Despite this initial delay, the bar eventually resumes steady growth at a later stage (bottom row).
@@
Tidal perturbations can also postpone the onset of bar formation. Figure 6 illustrates a delayed scenario in the cold disk model ($M_{\mathrm{per}}/M_{\mathrm{gal}}=1/3$, perpendicular orbit). The first row shows a moment shortly before the perturber’s closest approach. Although weak, the cold disk has already spontaneously generated a “seed bar” prior to the interaction. When the perturbation wave reaches the galaxy center, it interferes destructively with the original seed bar, which twists the quadrupole signal in the velocity fields of the central region and disrupts the spiral arms in the outskirts. The destructive interference limits the bar’s growth. Ultimately, one mode prevails over the other at a later time and starts the steady growth of the bar as seen in the bottom row ($A_{2}=0.1$). We propose that the internal mode dominates over the external mode as the perturber is already more than $300\;R_{d}$ away when $A_{2}$ reaches 0.1.
@@
Our results align with the interpretation of Moetazedian et al. (2017) regarding the delayed case. Destructive interference is due to the external perturbation wave being accidentally out of phase with the internally developed seed bar when it reaches the galaxy center. We test this hypothesis by resimulating the delayed case with slightly altered pericenter times, shifted by ±$\pm 0.1$ Gyr. In both the earlier and later interactions, the onset of bar formation is advanced by approximately 0.2 Gyr compared to the isolated cases, rather than being delayed as in the original interaction. This result underscores the difficulty for tidal perturbations to delay bar formation within galaxies inherently susceptible to internal instability, thereby explaining why the majority of our interaction simulations have earlier bar emergences. Tidal perturbations provide stronger initial noises for the swing amplification feedback loop, while the probability of the external perturbation wave being out of phase with the internally developed seed bar is indeed small. Our results also highlight the complexity of predicting the galaxy morphology, which is influenced by both their environment and internal baryonic physics (Zhou et al., 2020).
4.3 bar growth in the hot disk model
@@
As for bars in galaxies that are stable against bar formation when isolated (“tidally forced bars”), it is unclear whether their growth timescales are similar regardless of the perturbation strength. Consequently, we examined bar growth in the hot disk model, which avoids bar instability throughout a 6 Gyr isolated evolution.
@@
We utilize the same exponential fitting (Equation 4) for the $A_{2}$ evolution in the hot disk model. Since $A_{2}$ does not exhibit a distinct peak in the hot disk, we extend the upper limit of the fitting time range to the end of the simulation, i.e., 6 Gyr. For the lower limit, we set $A_{2}>0.025$ and exclude the initial 0.75 Gyr of the simulation. Visual examination of the stellar surface density indicates that the significant increase in the $A_{2}$ signal around the interaction time ($t_{\rm peri}=0.5$ Gyr) is primarily due to overall distortion of the disk and transient spirals, rather than bar formation. The fitting outcomes are presented in Figure 7. Unlike the cold and warm disks, the hot disk does not display a consistent $\tau_{\mathrm{bar}}$, ranging from a minimum of 1.85 Gyr to a maximum of 6.11 Gyr. Notably, the $M_{\mathrm{per}}/M_{\mathrm{gal}}=1/1$ interactions yield larger $\tau_{\mathrm{bar}}$values compared to the interactions with less massive perturbers, even though the bars in the former cases are stronger. When $M_{\mathrm{per}}/M_{\mathrm{gal}}$ is fixed, the $\tau_{\mathrm{bar}}$ values are larger for bars developed under prograde interactions compared to those that form under perpendicular and retrograde interactions. Since perturbers on prograde orbit provide stronger perturbations compared to those in perpendicular or retrograde orbit (see Łokas 2018 and Sec 4.2 of Paper I), this result also implies that stronger perturbations result in larger $\tau_{\mathrm{bar}}$ values. We suggest that the $\tau_{\mathrm{bar}}$ of the tidally forced bar qualitatively correlates with the strength of the perturbation, and leave quantification of the correlation for further investigation in the future.  [^3] One way to measure the perturbation strength by the dimensionless tidal strength parameter $S$ introduced in Elmegreen et al. (1991). We found that $S$ is approximately 0.87, 0.15, and 0.03 for $M_{\mathrm{per}}/M_{\mathrm{gal}}=1/1$, 1/3, and 1/10 respectively. However, $S$ does not take into account the inclination of the perturber’s orbit that also affects the strength of tidal interaction (see Łokas 2018; Sec 4.2 of Paper I).
@@
where $k$ represents the linear growth rate of the bar strength. The linear fitting is also presented in Figure 7 with the $k$ values noted in the text box. Similar to the result for $\tau_{\mathrm{bar}}$, we find a qualitative relation that stronger interactions correspond to larger $k$ values, suggesting a correlation between interaction strength and bar growth in the hot disk.
Visually, the linear function seems to offer a slightly better fit to the growth of tidally forced bars in the hot disk for most scenarios. From a statistical perspective, the linear fit yields a smaller reduced chi-squared statistic ($\chi_{\nu}^{2}$) compared to the exponential fit. Conversely, the cold and warm disks exhibit a more favorable fit with the exponential function during the bar formation phase. The linear fitting approach may be more appropriate for describing the secular evolution of the bar in the cold and warm disks, occurring after the buckling stage and prior to the saturation stage (refer to the inset axes in Figure 1, and also Figure 4 in Paper I).
The linear growth can be interpreted as an approximation of exponential growth with a large timescale. If a series of models of increasing stability is simulated, it is expected that the increasing $\tau_{\mathrm{bar}}$ will gradually transition the bar growth from exponential to linear. Therefore, the slight preference for linear bar growth alone is insufficient to serve as decisive evidence that tidally forced bars in the hot disk model are fundamentally different from those developed in the cold and warm disks.
However, we emphasize an important fact that both the exponential and linear growth rates of the tidally forced bars in the hot disk correlate with the perturbation strength. This contrasts with the consistent growth rate of bars within cold or warm disks, where bar growth is primarily driven by internal disk instability. This difference makes it less meaningful to compare the growth timescale $\tau_{\mathrm{bar}}$ between the hot disk and the cold/warm disks. It also suggests that bars in the hot disk may not operate under the swing amplification mechanism, contrasting with those in the cold and warm disks. Possibly, the hot dynamical state of the disk resists the disturbance caused by the perturber, preventing disruption of the galaxy center and rendering it less susceptible to the swing amplification feedback loop (Dubinski et al., 2008).
@@
The formation and evolution of tidally forced bars differ from those in galaxies that are prone to bar instability. Quantitative comparison of bar shapes and dynamical/kinematical analysis of the bar regions have the potential to distinguish these two categories of bars. This will constitute the long-term objective of this series of papers.
@@ 5 Summary
We apply an exponential fitting to the $A_{2}$ evolution in cold and warm disks to quantify the growth timescale $\tau_{\mathrm{bar}}$ and the onset time $t_{0}$ of the bar formation (Figure 1). The tidally induced bars show similar $\tau_{\mathrm{bar}}$ to that of the internally developed bars with differences less than 20%, indicating that tidal bars have similar growth rates to their spontaneous counterpart in the same disk (Figure 2). The tidal perturbation promotes(delays) the bar formation by advancing(postponing) the onset time $t_{0}$. This conclusion is further corroborated by comparing the $A_{2}$ evolution against $(t-t_{0})$. (Figure 3). If the disk inherently possesses instability, the primary driver of bar formation is its internal nature, which remains largely unaffected by tidal perturbations.
We further investigate the onset of the bar formation by examining the perturbation wave. In instances where the perturbation advances bar formation, the perturbation wave propagates to the galaxy center shortly after the perturber’s pericenter and evokes a “seed bar” (Figure 4). The isolated case shows a similar evolution of the “seed bar” but its spontaneous emergence occurs approximately 0.7 Gyr later than in the interaction case (Figure 5). We also show a delayed interaction case where the perturbation wave is out of phase with the internally developed seed bar, leading to destructive interference and limiting the growth of the bar (Figure 6).
In the hot disk model, the tidally forced bars do not show consistent $\tau_{\mathrm{bar}}$ values (Figure 7), probably due to deviations from exponential growth profiles. Linear fitting provides a slightly better description of the bar growth in the hot disk than the exponential fitting. These results make it less meaningful to compare the growth timescale $\tau_{\mathrm{bar}}$ between the hot disk and the cold/warm disks and hints that bars in the hot disk may not follow the same growth mechanism as those in the cold and warm disks.
However, the difference in the bar growth mechanisms holds the potential to separate tidally forced bars that developed in stable galaxies from those developed in galaxies that are inherently susceptible to bar instability. Further investigation of the bar’s shape and dynamical/kinematical analysis of the bar regions could be important in distinguishing between these two types of bars.
With the advent of powerful telescopes, such as the Atacama Large Millimeter Array and the James Webb Space Telescope, astronomers can now observe the gas-rich phase of galaxies at high redshifts. Considering that all disk galaxies undergo a gas-rich phase, the new era of gas-rich controlled simulations is vital for further insights into the bar formation mechanism (Bland-Hawthorn et al., 2024, 2025). We will explore the impact of gas on bar formation and evolution under both the internal and external formation mechanisms in future work.
@@ Acknowledgements
We thank the referee for suggestions that helped to improve the presentation of the paper. We thank Thor Tepper-Garcia and Joss Bland-Hawthorn for their help with the agama work and with initializing the galaxy models. We thank Zhi Li and Sandeep Kumar Kataria for their valuable insights on simulations and analysis. We also thank Rui Guo for their helpful discussions. The research presented here is partially supported by the National Key R&D Program of China under grant No. 2018YFA0404501; by the National Natural Science Foundation of China under grant Nos. 12025302, 11773052, and 11761131016; by the “111” Project of the Ministry of Education of China under grant No. B20019; and by China Manned Space Program with grant No. CMS-CSST-2025-A11. J.S. acknowledges support from the Newton Advanced Fellowship awarded by the Royal Society and the Newton Fund. X.W. wishes to thank the financial support from the Natural Science Foundation of China (Numbers NSFC-12073026 and NSFC-12433002). This work made use of the Gravity Supercomputer at the Department of Astronomy, Shanghai Jiao Tong University.